Solve: dy/dx + 2xy= e^-x^2
1
Expert's answer
2020-09-15T16:30:31-0400

"\displaystyle\frac{\mathrm{d}y}{\mathrm{d}x} + 2xy = e^{-x^2} \\



\textsf{For the ODE}\hspace{0.1cm} \frac{\mathrm{d}y}{\mathrm{d}x} + Py = Q, \\ 



\textsf{Where P and Q are functions} \\\textsf{of x, it has the solution} \\



y IF = \int Q\cdot IF\hspace{0.1cm} \mathrm{d}x, \\



\textsf{Where IF is the integrating}\\\textsf{factor}\\



IF = e^{\int P \hspace{0.1cm} \mathrm{d}x} = e^{\int 2x \hspace{0.1cm} \mathrm{d}x} = e^{x^2} \\



\therefore ye^{x^2} = \int e^{-x^2} \cdot e^{x^2} \hspace{0.1cm} \mathrm{d}x\\ = \int e^0 \mathrm{d}x = \int \mathrm{d}x = x + C \\



\therefore ye^{x^2} = x + C, \Rightarrow\\ y = e^{-x^2}(x + C) \\ \textit{Where C is an arbitrary constant}\\



\therefore y = e^{-x^2}(x + C)\\ \textsf{is the solution to the differential}\\\textsf{equation}"